OpenStudy (anonymous):

Find the distance between the pts. and find the coordinates of the midpoint of the segment with the given pts. R(5,2) S(-2,4)

OpenStudy (anonymous):

So, we start by using Pythagorean's theorem to determine the magnitude of the distance between the two points: \[Distance = \sqrt{(5-(-2))^{2}+(2-4)^{2}}\] \[Distance =\sqrt{53}\] Then we find the midpoint between the points by taking an "Average" of the the sum of both x coordinates in S and R, and the Average of the y coordinates in S and R: \[Midpoint = (\frac{ 5+(-2) }{ 2 }, \frac{2+4}{2})\] \[Midpoint = (\frac{3}{2}, 3)\]